OpenStudy (pogo):

1st person to comment gets a chance to kiss me :3

OpenStudy (christos):

1+v+v^2 = (1+v)^2

OpenStudy (christos):

1^2 + 2v + v^2 = 1+v+v^2

OpenStudy (christos):

solve for v